5) At a work rate of 300 Watts, Puja’s heart rate was 150 bpm, exhaled %O2 was 17%, exhaled %CO2 was 5%,

and minute volume was 40L/Min. Calculate Puja’s RQ.

A) 0.5

B) 0.8

C) 1.0

D) 1.25---- ANSWER
Answer #1

5.% of oxygen in atmosphere = 21%

% of oxygen given out = 17%

So % of oxygen taken in = 4%

Total amount of oxygen taken in = 4×40/100 = 1.6L/min.

% of CO2 released out = 5%

Total amount of CO2 = 5× 40/100=2L/min.

RQ = amount of CO2 released/amountvof CO2 taken in = 5/4 = 1.25